A pair of narrow slits is illuminated with light of wavelength λ= 542.1 nm. The resulting interference maxima are found to be separated by 0.92 mm on a screen 0.86 m from the slits. What is the separation of the slits? (mm)

Please show your work!
Answer #1

Seperation between interference manna = els d ß = 20 - d = 20 , 7=542.1x107m D=0.86m ß = 0.92 x 103 m - d = 542-1x109x0096 0.